# Collisional whistler instability and electron temperature staircase in inhomogeneous plasma

### v1

@@ 1 Introduction
X-ray observations of the diffuse plasma residing within galaxy clusters have revealed intricately structured temperature fields whose sharp gradients are inferred to have persisted far longer than classical transport theory predicts (Peterson & Fabian, 2006; Markevitch & Vikhlinin, 2007). Recent experiments on NIF with laser-produced hot, magnetized, high-beta plasmas also feature such anomalously structured temperature fields (Meinecke et al., 2022). Such structure would require a significantly reduced level of electron heat transport, which might be due to plasma microinstabilities. Indeed, it is now well-established (Kunz et al., 2022) that the macroscopic transport properties of high-beta magnetized plasmas are significantly modified by small-scale instabilities such as firehose (Rosin et al., 2011; Kunz et al., 2014), mirror (Kunz et al., 2014; Komarov et al., 2016), or heat-flux whistler instabilities (Levinson & Eichler, 1992; Pistinner & Eichler, 1998; Komarov et al., 2018; Roberg-Clark et al., 2018; Drake et al., 2021).
@@
Here we remove this shortcoming by deriving the Wigner–Moyal equations that govern the collisional whistler instability in a slab geometry, similar to what has been done in modelling drift-wave turbulence beyond the geometrical-optics approximation (Ruiz et al., 2016; Zhu et al., 2018; Tsiolis et al., 2020). We find additional terms in the instability dispersion relation and growth rate due to gradients in the background plasma. We proceed to show that these additional terms can actually stabilize a non-constant temperature profile along the magnetic field, which is not possible if only the geometrical-optics approximation is used. Equivalently, these additional terms cause the instability to be damped at low temperatures, providing a mechanism for the instability to re-arrange the temperature profile into a marginally stable state via excitation at high temperature, propagation down the temperature gradient via Nernst advection, and subsequent damping at low temperature. We proceed to show that the marginally stable temperature profile generically takes the form of a staircase where isothermal regions are insulated from each other by abrupt jumps in the temperature, which occur at the zeros of a certain function comprised of an intricate combination of magnetic transport coefficients. These staircases can be in pressure balance, resembling the ubiquitous cold fronts in galaxy clusters (Markevitch & Vikhlinin, 2007) but with the magnetic field no longer required to drape the front.
@@
This paper is organized as follows. In Sec. 2, the governing equations and slab geometry are introduced, and some initial simplifications are performed. In Sec. 3, the dispersion relation and growth rate of the collisional whistler instability is obtained. In Sec. 4, the stability criterion is derived, and the underlying physics of the additional stabilization terms due to background gradients is discussed. In Sec. 5, the instability growth rate is compared with the conduction time of the background temperature profile to determine the parameter regime in which the instability will be dynamically relevant. In Sec. 6, the marginally stable temperature profile is derived and shown generically to have a staircase structure. Isobaric and constant-density plasmas are discussed explicitly as special cases. In Sec. 7, the back-reaction of the instability on the background temperature profile is derived and regimes in which the work done by friction is a net cooling effect and also when the heat flux can be reduced by the instability are outlined. In Sec. 8, the main results are summarized. Auxiliary calculations are presented in appendices.
@@ 2.1 Extended electron MHD equations in slab geometry
×$\nabla\times{\boldsymbol{\rm B}}(t,z)=\begin{pmatrix}-\partial_{z}\widetilde{B}_{y}(t,z)\\ \partial_{z}\widetilde{B}_{x}(t,z)\\ 0\end{pmatrix}=\begin{pmatrix}-\mathsf{J}&0\\ 0&0\end{pmatrix}\partial_{z}{\boldsymbol{\rm B}},$ (4)
@@ 2.3 Chapman–Enskog friction coefficients
Equations (12) and (13) are valid for any friction force, allowing one to study driven systems. For undriven systems, the friction force is determined by the plasma fluid variables themselves according to some closure. A common closure that we shall adopt here is provided by the Chapman–Enskog method (Helander & Sigmar, 2002; Bott et al., 2024), which for the Lorentz collision operator yields the following expressions for the friction force and for the heat flux (Epperlein, 1984):
@@
where $\tau_{ei}$ is the electron-ion collision time (Helander & Sigmar, 2002; Epperlein & Haines, 1986) and $v_{t}$ is the thermal speed, defined respectively as
@@
In the remainder of the analysis, we shall use the rational interpolants for the various transport coefficients as functions of $\mathcal{M}$ provided by Davies et al. (2021) with the asymptotic corrections for the Lorentz operator developed in Lopez (2024). Their limiting forms as $\mathcal{M}\to 0$ and $\mathcal{M}\to\infty$ are listed in Appendix B.
@@ 2.4 Resulting equations for Chapman–Enskog friction forces
The Chapman–Enskog heat flux (14) can be shown to take the form
@@
$\frac{3}{2}n\,\partial_{t}T$ $=\epsilon^{2}\frac{B_{z}^{2}}{8\pi}\left[\eta|\partial_{z}\psi|^{2}-u_{\gamma}\textrm{Im}\left(\psi^{*}\partial_{z}\psi\right)-\frac{v_{N}}{2}\partial_{z}|\psi|^{2}\right]-\partial_{z}q_{z}.$ (26)
@@ 3 Dispersion relation, growth rate, and breakdown of the short-wavelength approximation
where $\mathcal{M}=\Omega\tau_{ei}$. In this limit, the physical interpretation of these quantities becomes clearer: $G_{d}$ represents the familiar group-velocity dispersion for whistler waves but modified by friction-induced Hall effect (Davies et al., 2021), $\eta$ governs the resistive diffusion of magnetic-field perturbations, $u_{\gamma}$ is the cross-gradient Nernst advection velocity, and $v_{N}$ is the standard Nernst advection velocity.
@@ 3.1 Derivation via Wigner–Moyal phase-space formulation
In our first approach, we shall derive the phase-space analog of (22) that governs the Wigner function of the complex mode amplitude $\psi$, defined as
@@
This approach will bring the Hamiltonian structure of (22) to light, allowing us then to extract the dispersion relation and growth rate immediately.
@@
$\smash{\widehat{D}}=\smash{\widehat{k}}_{z}\left[G_{d}(\smash{\widehat{z}})-i\eta(\smash{\widehat{z}})\right]\smash{\widehat{k}}_{z}+\smash{\widehat{k}}_{z}\left[v_{N}(\smash{\widehat{z}})+iu_{\gamma}(\smash{\widehat{z}})\right].$ (32)
@@
Finally, applying the Wigner–Weyl transform (WWT) [^1] For a review of the WWT, see Appendix C. yields the Wigner–Moyal kinetic equation that governs the Wigner function (30) of the fluctuations:
@@
3.2 Derivation via single-mode polar decomposition
As an alternative means of deriving the Hamiltonian (35), let us instead represent $\psi$ by its polar decomposition
$\psi=\sqrt{\mathcal{I}}\,\exp(i\theta).$ (38)
Thus, we see that the Hermitian and anti-Hermitian parts of the Hamiltonian identified in (35) via WWT-based methods emerges from the traditional approach as well. In fact, the evolution equation for the polar amplitude (41) is identical to that for the wave intensity given by (36). This is because, as discussed further in Appendix D, the final set of gradient terms in (41) encode the ‘non-eikonal’ bandwidth $\left\langle k_{z}^{2}\right\rangle-\left\langle k_{z}\right\rangle^{2}$ that can be combined with the term $\mathcal{D}_{A}\left(\partial_{z}\theta,z\right)$ to yield the Wigner-averaged growth rate $\left\langle\mathcal{D}_{A}\right\rangle$. The non-eikonal bandwidth being automatically contained in the Wigner-based formalism, rather than being a separate term that must be included in the evolution equations, is a theoretical advantage of that approach.
@@ 3.3 Insufficiency of short-wavelength approximation
Since $u_{\gamma}\propto\partial_{z}T$, (42) shows that the fluctuation wavelength may be comparable to, or even smaller than, the temperature lengthscale $L_{T}=(\partial_{z}\log T)^{-1}$, depending on the prefactor. Indeed, one has
@@
Let us conclude this section with a brief discussion regarding the relevance of our analysis to current laser-plasma experiments. Consider the initial stages of an experiment such as that performed by Meinecke et al. (2022). In such an experiment, pressure balance is quickly established before self-generated magnetic fields have time to grow to appreciable strength (there are no imposed zeroth-order fields). Thus, we can view the initial phase of the experiment as residing within the upper-left corner of Fig. 5 (low $\mathcal{M}$ and high $\beta_{0}$). As time progresses, dynamo action causes magnetic fields to grow while maintaining constant pressure, so the system evolves to a higher $\mathcal{M}$ state along the trajectory $\beta_{0}\sim\mathcal{M}^{-2}$ (if temperature is not constant during this time, then the evolution of $\mathcal{M}$ is even faster, following the shallower trajectory $\beta_{0}\sim T^{5}\mathcal{M}^{-2}$). Along such a trajectory, $k_{z,\text{max}}L$ will be an increasing function since the contours go as $\beta_{0}\sim\mathcal{M}^{-3}$ (43). There is a subtlety, however, in that the maximum growth rate for the collisional whistler instability is negative when $k_{z,\text{max}}L\lesssim 1$, as will be discussed later (see Sec. 4 and also Fig. 9). This means that whistler waves will not be excited in the experiment until the plasma is magnetized enough so that $k_{z,\text{max}}L\sim 1$, at which point modes whose wavelengths are comparable with the gradient lengthscale will appear. Due to their early excitation, these modes will have the most time subsequently to manipulate the plasma evolution (see Sec. 7), but due to their long wavelengths, they can only be accurately described by the Wigner–Moyal analysis performed here.

### v2

@@ 1 Introduction
X-ray observations of the diffuse plasma residing within galaxy clusters have revealed intricately structured temperature fields whose sharp gradients are inferred to have persisted far longer than classical transport theory predicts (Peterson and Fabian, 2006; Markevitch and Vikhlinin, 2007). Recent experiments on NIF with laser-produced hot, magnetized, high-beta plasmas also feature such anomalously structured temperature fields (Meinecke et al., 2022). Such structure would require a significantly reduced level of electron heat transport, which might be due to plasma microinstabilities. Indeed, it is now well-established (Kunz et al., 2022) that the macroscopic transport properties of high-beta magnetized plasmas are significantly modified by small-scale instabilities such as firehose (Rosin et al., 2011; Kunz et al., 2014), mirror (Kunz et al., 2014; Komarov et al., 2016), or heat-flux whistler instabilities (Levinson and Eichler, 1992; Pistinner and Eichler, 1998; Komarov et al., 2018; Roberg-Clark et al., 2018; Drake et al., 2021; Yerger et al., 2024).
@@
Here we remove this shortcoming by deriving the Wigner–Moyal equations that govern the collisional whistler instability in a slab geometry, similar to what has been done in modelling drift-wave turbulence beyond the geometrical-optics approximation (Ruiz et al., 2016; Zhu et al., 2018; Tsiolis et al., 2020). We find additional terms in the instability dispersion relation and growth rate due to gradients in the background plasma. We proceed to show that these additional terms can actually stabilize a non-constant temperature profile along the magnetic field, which is not possible if only the geometrical-optics approximation is used. Equivalently, these additional terms cause the instability to be damped at low temperatures, providing a mechanism for the instability to re-arrange the temperature profile into a marginally stable state via excitation at high temperature, propagation down the temperature gradient via Nernst advection, and subsequent damping at low temperature. We proceed to show that the marginally stable temperature profile generically takes the form of a staircase where isothermal regions are insulated from each other by abrupt jumps in the temperature, which occur at the zeros of a certain function comprised of an intricate combination of magnetic transport coefficients. These staircases can be in pressure balance, resembling the ubiquitous cold fronts in galaxy clusters (Markevitch and Vikhlinin, 2007) but with the magnetic field no longer required to drape the front.
@@
2 Summary
Here we first provide an executive summary that highlights the main definitions, discussions, and results for each section, serving as an overall roadmap of the paper that can be consulted later for easy reference.
In Sec. 3, the governing electron MHD equations and slab geometry are introduced, leading to the set (6). Fundamentally, the electron-MHD limit in slab geometry results in the simplification that only the electron temperature and the perpendicular components of the magnetic field (with respect to the single direction of inhomogeneity) have non-trivial time evolution; the electron density and the parallel magnetic field both remain constant in time. The perpendicular magnetic field is then expressed in the diagonalizing eigenbasis for its evolution equation, leading to the simplified description in terms of mode amplitudes given in (12) and (13). This is all for a general, unspecified friction and heat flux; the remaining parts of this section (and the remainder of the paper) specialize to when the friction and heat flux are determined by the standard Chapman–Enskog expressions (14). The nine transport coefficients ($\alpha_{\parallel}$, $\alpha_{\perp}$, $\alpha_{\wedge}$, $\beta_{\parallel}$, $\beta_{\perp}$, $\beta_{\wedge}$, $\kappa_{\parallel}$, $\kappa_{\perp}$, $\kappa_{\wedge}$) are all generally functions of the dimensionless magnetization parameter $\mathcal{M}$ defined in (17). Ultimately, after inserting the same eigenmode decomposition into the expressions for the friction and heat flux, one arrives at the final set of working equations: the magnetic-field mode amplitudes are governed by (22) and the temperature evolution is governed by (26). Up to that point, all manipulations of the initial equations are exact, and all nonlinearities are retained.
In Sec. 4, the dynamical equations are linearized about the small transverse field amplitude in order to obtain their dispersion relation and growth rate; since the temperature evolution involves terms quadratic or more in the mode amplitude, the two dynamical equations decouple in this limit. The temperature profiles are treated as fixed in time from the perspective of the magnetic field fluctuations, even though heat conduction is still present in the lowest-order temperature evolution equation; the validity of this approximation is outlined in Sec. 6. To maintain an exact treatment (within the linear regime), the dispersion relation and growth rate are obtained as the Hamiltonian and non-Hamiltonian components of the generalized wave-kinetic equation (34) that governs the magnetic fluctuations. Specifically, the Hermitian and anti-Hermitian parts of the dispersion relation are given in (35), with the constituent terms defined in (27). This method of deriving the dispersion relation is chosen because it does not rely on making a short-wavelength approximation. The remainder of Sec. 4 is therefore dedicated to defining the parameter space in which such a general treatment in necessary; it is found that the wavelength for the maximally growing mode becomes comparable to the background medium inhomogeneity lengthscale in the regime $\mathcal{M}^{3}\beta_{0}\ll 1$, where $\beta_{0}$ is the plasma beta defined in (40).
In Sec. 5, the expression for the growth rate (35b) is examined in more detail to understand the condition for instability (42). Allowing for wavelengths comparable to inhomogeneity scale of the medium introduces two additional stabilizing mechanisms: one related to the gradient of the Nernst velocity that advects the perturbations, and one to the curvature of the plasma resistivity that diffuses the perturbations. Simple heuristic descriptions of these two stabilizing mechanisms are then presented briefly. The section concludes with a specific discussion for the simple case (45) when the plasma is in pressure balance with a linear temperature profile; in this case, it is shown that the additional stabilizing terms cause a transition from unstable to stable behaviour analogous to the transition of the short-wavelength approximation from being valid to being violated discussed in Sec. 4.
In Sec. 6, the validity for the approximation that the temperature profile remain fixed in time for the linear stability analysis is assessed by comparing the instability growth rate (46) with the diffusion time (49) of the background temperature profile. This analysis accommodates arbitrarily shaped plasma profiles, with local lengthscales and curvatures defined in (47). For convenience, the growth rate (46) is also split into terms driven by temperature gradients, temperature curvatures, mixed temperature and density gradients, density gradients, and density curvatures; the coefficient functions, defined in (48), are grouped together respectively as $F_{T,1}$, $F_{T,2}$, $F_{T,n}$, $F_{n,1}$, and $F_{n,2}$. Generally, it is found that the instability’s growth rate (50) normalised by the dynamical evolution time is only greater than unity in a sub-region of parameter space where both $\mathcal{M}$ and $\beta_{0}$ are large; this is in contrast to the short-wavelength approximation that predicts the instability growth to be dynamically relevant everywhere [cf. (54)]. However, this result is nuanced because only regions where the instability is present are considered, and in fact, in much of the excluded parameter space there is no instability and perturbations are instead strongly damped (at a rate much larger than the diffusion time). The strong damping is due to the additional stabilizing mechanisms that occur when the wavelengths become comparable to the inhomogeneity scale discussed in Sec. 5.
In Sec. 7, the marginally stable temperature profile (satisfying the condition for zero growth rate everywhere) is derived. Due to the additional stabilizing terms in the expression for the growth rate, the solution involves non-trivial global structure. As discussed first in the general case, the governing equation (63) for the marginally stable $\mathcal{M}$ profile (a proxy for temperature) has the general structure of a nonlinear boundary-layer differential equation (67), due to the singular nature as $\beta_{0}$ becomes large of the coefficient $g(\mathcal{M})$ defined in (64) and (60). A variable transformation is performed to recast the nonlinear differential equation into a linear one that is then readily solved by successive integrations, yielding the solution (66) for the inverse function of the spatial profile $\mathcal{M}(z)$. This solution is then shown to exhibit a staircase structure as a generic feature due to the boundary layers. The remainder of the section specializes the general theory to two situations: plasmas with constant pressure and plasmas with constant density. Both cases are qualitatively similar when viewed in a parameter space consisting of $\mathcal{M}$ and some measure of the plasma beta; for the constant-pressure case, this is simply $\beta_{0}$, but for the constant density case, an effective plasma beta is defined in (84). The staircase is shown to have a single step, in rough correspondence to the transition between (dynamically relevant) instability growth and damping outlined in the previous section 6. The predicted staircase feature of the solution is then verified with direct numerical simulation of the governing differential equation along with a closed-form analytical approximation described in (78).
In Sec. 8, the linear analysis is extended to a quasilinear study of the electron temperature response by restoring the lowest-order (quadratic) fluctuation terms in the temperature-evolution equation (86). Two different groupings of the various terms are presented depending on the physics to be emphasized, with definitions provided in (87). First, it is discussed how (as required by energy conservation), the growth of unstable perturbations also corresponds to a net cooling effect of the friction forces when the condition (91) is satisfied. Next, the total heat flux (including Ettingshausen terms generated by the instability) is analyzed, and found to be reduced compared to the standard conductive heat flux when the condition (94) is satisfied, although the reduction is generally modest. Lastly, the heat-flux reduction is analyzed when the temperature has the marginally stable staircase profile derived in the previous section 7. The extreme situation in which the heat flux is completely suppressed for the marginal-stability profiles is then considered to derive required conditions on the instability intensity profiles; the resulting expression (100) suggests that such high degrees of heat-flux suppression are unlikely to occur for high-beta plasmas within the validity of the quasilinear fluid model adopted here.
Finally, in Sec. 9, the main results are summarized and directions for future work are outlined. Auxiliary calculations and review material are presented in appendices.
@@ 3.1 Extended electron MHD equations in slab geometry
$\nabla\times{\boldsymbol{\rm B}}(t,z)=\begin{pmatrix}-\partial_{z}\widetilde{B}_{y}(t,z)\\ \partial_{z}\widetilde{B}_{x}(t,z)\\ 0\end{pmatrix}=\begin{pmatrix}-\mathsf{J}&0\\ 0&0\end{pmatrix}\partial_{z}{\boldsymbol{\rm B}},$ (4)
@@ 3.3 Chapman–Enskog friction coefficients
Equations (12) and (13) are valid for any friction force, allowing one to study driven systems. For undriven systems, the friction force is determined by the plasma fluid variables themselves according to some closure. A common closure that we shall adopt here is provided by the Chapman–Enskog method (Helander and Sigmar, 2002; Bott et al., 2024), which for the Lorentz collision operator yields the following expressions for the friction force and for the heat flux (Epperlein, 1984):
@@
where $\tau_{ei}$ is the electron-ion collision time (Helander and Sigmar, 2002; Epperlein and Haines, 1986) and $v_{t}$ is the thermal speed, defined respectively as
@@
In the remainder of the analysis, we shall use the rational interpolants for the various transport coefficients as functions of $\mathcal{M}$ developed in Lopez (2024). Their limiting forms as $\mathcal{M}\to 0$ and $\mathcal{M}\to\infty$ are listed in Appendix B.
@@ 3.4 Resulting equations for Chapman–Enskog friction forces
The $z$-component of Chapman–Enskog heat flux (14), $q_{z}$, that appears in (13) can be shown to take the form
@@
$\frac{3}{2}n\,\partial_{t}T$ $=\epsilon^{2}\frac{B_{z}^{2}}{8\pi}\left[\eta|\partial_{z}\psi|^{2}-u_{\gamma}\textrm{Im}\left(\psi^{*}\partial_{z}\psi\right)-\frac{v_{N}}{2}\partial_{z}|\psi|^{2}\right]-\partial_{z}q_{z},$ (26)
with $q_{z}$ given in (24).
In summary, the two main equations we shall be working with in the remainder of this paper are the evolution of the transverse field perturbations (describing the linear instability dynamics) governed by (22), and the evolution of the temperature profile (describing the back-reaction of the instability on the equilibrium dynamics) governed by (26). More will be said about the physical content of (26) in Sec. 8, but as a brief preview, it is worthwhile to highlight now that only the first term within square brackets is manifestly positive (corresponding to resistive heating). The remaining terms, although still related to frictional forces, can take either sign depending on whether the instability is growing or damping, allowing the instability to re-distribute the temperature profile. As we shall discuss in later sections (beginning in the following section), this effect can be particularly prominent in the high-beta, low-magnetization regime.
@@ 4 Dispersion relation, growth rate, and breakdown of the short-wavelength approximation
where $\mathcal{M}=\Omega\tau_{ei}$. In this limit, the physical interpretation of these quantities becomes clearer: $G_{d}$ represents the familiar group-velocity dispersion for whistler waves but modified by friction-induced Hall effect (Davies et al., 2021), $\eta$ governs the resistive diffusion of magnetic-field perturbations, $u_{\gamma}$ is the cross-gradient Nernst advection velocity, and $v_{N}$ is the standard Nernst advection velocity. In particular, the presence of $G_{d}$ explains our choice to call this instability the ‘collisional whistler instability’; in the absence of friction terms (i.e., with all $\alpha$ and $\beta$ coefficients set to zero), the dispersion relation (35a), which we shall derive shortly, would be identical to the standard whistler dispersion relation in the electron-MHD limit [see, for example, Komarov et al. (2018)].
@@ 4.1 Derivation via Wigner–Moyal phase-space formulation
Here we shall derive the phase-space analog of (22) that governs the Wigner function of the complex mode amplitude $\psi$, defined as
@@
This will bring the Hamiltonian structure of (22) to light, allowing us then to extract the dispersion relation and growth rate immediately. The following derivation closely follows the presentation of Ruiz et al. (2016), so the reader is invited to consult that reference along with the brief summary provided in Appendix C for more detailed explanations of the steps involved. A more conventional derivation of the same dispersion relation and growth rate, based on a polar decomposition of the wavefield into an amplitude and phase, is presented in Appendix D.
@@
[…]
@@
Finally, applying the Wigner–Weyl transform (WWT, see Appendix C) yields the Wigner–Moyal kinetic equation that governs the Wigner function (30) of the fluctuations:
@@ 4.2 Insufficiency of short-wavelength approximation
Since $u_{\gamma}\propto\partial_{z}T$, (38) shows that the fluctuation wavelength may be comparable to, or even larger than, the temperature lengthscale $L_{T}=(\partial_{z}\log T)^{-1}$, depending on the prefactor. Indeed, one has
@@
Let us conclude this section with a brief discussion regarding the relevance of our analysis to current laser-plasma experiments. Consider the initial stages of an experiment such as that performed by Meinecke et al. (2022). In such an experiment, pressure balance is quickly established before self-generated magnetic fields have time to grow to appreciable strength (there are no imposed zeroth-order fields). Thus, we can view the initial phase of the experiment as residing within the upper-left corner of Fig. 5 (low $\mathcal{M}$ and high $\beta_{0}$). As time progresses, dynamo action causes magnetic fields to grow while maintaining constant pressure, so the system evolves to a higher $\mathcal{M}$ state along the trajectory $\beta_{0}\sim\mathcal{M}^{-2}$ (if temperature is not constant during this time, then the evolution of $\mathcal{M}$ is even faster, following the shallower trajectory $\beta_{0}\sim T^{5}\mathcal{M}^{-2}$). Along such a trajectory, $k_{z,\text{max}}L$ will be an increasing function since the contours go as $\beta_{0}\sim\mathcal{M}^{-3}$ (39). There is a subtlety, however, in that the maximum growth rate for the collisional whistler instability is negative when $k_{z,\text{max}}L\lesssim 1$, as will be discussed later (see Sec. 5 and also Fig. 9). This means that whistler waves will not be excited in the experiment until the plasma is magnetized enough so that $k_{z,\text{max}}L\sim 1$, at which point modes whose wavelengths are comparable with the gradient lengthscale will appear. Due to their early excitation, these modes will have the most time subsequently to manipulate the plasma evolution [^1] This assumes the experiment progresses slowly enough for the time difference to be dynamically meaningful. (see Sec. 8), but due to their long wavelengths, they can only be accurately described by the Wigner–Moyal analysis performed here.